OpenStudy (anonymous):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –4 3 –2 4 Line 2 x y 0 8 2 6 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (anonymous):

@Crissy15 @demonchild99 @Great_And_Powerful @Ashleyrh1999 @horsegirl27 @Jennaluv

OpenStudy (demonchild99):

~facepalm~ i hate math....lol...or more or less it hates me...i dont know how to do this

OpenStudy (anonymous):

*high fives* thanx anyway

OpenStudy (horsegirl27):

I kind of understand what's happening but now really what I have to do, sorry. Plus I'm busy. But if you still need help later when I'm done I will try to help

OpenStudy (anonymous):

ok thank you @horsegirl27

OpenStudy (anonymous):

@k_lynn

OpenStudy (anonymous):

@monica10

OpenStudy (anonymous):

one solution

OpenStudy (anonymous):

@wingspansson.Silvera01

OpenStudy (anonymous):

@wingspansson.Silvera01

OpenStudy (anonymous):

It can only have 1 solution as the they are linear lines. Howevwr, if they are parallel, they will never cross and thus have no solution

OpenStudy (anonymous):

@jakashaka123 Next time, walk them through.

OpenStudy (anonymous):

im trying to but when i post what i need my post button disappears @Nibby

OpenStudy (anonymous):

line 1 x = -4 and -2 Y=3 and 4

OpenStudy (anonymous):

thank you all for the help

OpenStudy (anonymous):

slope = (4-3)/(-4--2) = 1/-2 = -1/2

OpenStudy (anonymous):

line 2 x=o and 2 and y = 8 and 6

OpenStudy (anonymous):

slope = (6-8)/(2-0) = -2/2 = -1

OpenStudy (anonymous):

Since the slopes are different, the line meet at one point. Ans:: One solution for the system

OpenStudy (anonymous):

ok thank you @jakashaka123

OpenStudy (anonymous):

srry for some reason my post button wasnt working

OpenStudy (anonymous):

and no problem @wingspansson.Silvera01